# Recurrent Aligned Network for Generalized Pedestrian Trajectory Prediction

### v1

@@ Abstract
Pedestrian trajectory prediction is a crucial component in computer vision and robotics, but remains challenging due to the domain shift problem. Previous studies have tried to tackle this problem by leveraging a portion of the trajectory data from the target domain to adapt the model. However, such domain adaptation methods are impractical in real-world scenarios, as it is infeasible to collect trajectory data from all potential target domains. In this paper, we study a task named generalized pedestrian trajectory prediction, with the aim of generalizing the model to unseen domains without accessing their trajectories. To tackle this task, we introduce a Recurrent Aligned Network (RAN) to minimize the domain gap through domain alignment. Specifically, we devise a recurrent alignment module to effectively align the trajectory feature spaces at both time-state and time-sequence levels by the recurrent alignment strategy. Furthermore, we introduce a pre-aligned representation module to combine social interactions with the recurrent alignment strategy, which aims to consider social interactions during the alignment process instead of just target trajectories. We extensively evaluate our method and compare it with state-of-the-art methods on three widely used benchmarks. The experimental results demonstrate the superior generalization capability of our method. Our work not only fills the gap in the generalization setting for practical pedestrian trajectory prediction but also sets strong baselines in this field.
Keywords:
Pedestrian Trajectory Prediction, Domain Generalization
1 Introduction
@@
Pedestrian trajectory prediction aims to estimate the future movements of pedestrians given their observed trajectories. This task is crucial in various applications such as autonomous driving [15, 21, 10], visual recognition [32, 7, 16], navigation [8, 12, 45], and video surveillance [28, 46, 6]. Despite significant progress in this field, pedestrian trajectory prediction remains challenging due to the domain shift problem. This problem arises when a model encounters scenes significantly different from its training dataset, resulting in performance degradation. For example, the motion styles of students on a university campus differ significantly from those of players on a basketball court. Consequently, a predictor trained on campus datasets (e.g., SDD) may excel in that scene, but encounter difficulties with the basketball court datasets (e.g., NBA), as depicted in Figure 1(a). Recent studies [41, 11] have identified the domain shift problem in pedestrian trajectory prediction and mitigate it through domain adaptation. However, in real-world scenarios like autonomous driving, domain changing is inevitable. Collecting trajectory data for all possible target domains is impractical, making the application of domain adaptation using data from target domains for model fine-tuning unfeasible in practice. In this paper, we tackle the challenge of predicting pedestrian trajectories in an unseen target domain dataset, without requiring data from the target domain during model training, as illustrated in Figure 1(b). We term this task as Generalized Pedestrian Trajectory Prediction. One effective method to address this challenge is through domain alignment, which minimizes the data distribution gap between the source and target domains in high-dimensional feature spaces by using alignment losses such as Maximum Mean Discrepancy (MMD) [22], Correlation Alignment Distance (CORAL) [33], and adversarial loss [39]. Nevertheless, traditional domain alignment strategies are not well suited for pedestrian trajectory prediction tasks because 1) they align different domains directly, without considering the alignment at both time-state and time-sequence levels within trajectories; 2) they overlook the surrounding alignment, ignoring that the predicted trajectory is not only associated with that specific trajectory but also related to other interconnected trajectories within the scene. To address these issues, we introduce a Recurrent Aligned Network (RAN), a new alignment framework tailored for generalized trajectory prediction, as illustrated in Figure 2. RAN deals with domain alignment by considering both time-state and time-sequence alignments, along with the social interaction dynamics inherent in pedestrian trajectories. This advancement is facilitated by two key modules: a recurrent alignment module and a pre-aligned representation module. The recurrent alignment module aims to minimize the domain gap at both time-state and time-sequence levels by a novel recurrent alignment strategy. Specifically, we first input trajectory representations from different domains at the starting timestamp into the GRU units and update the network using alignment losses calculated by the output latent variables. Then, we input the trajectory representations of the next timestamp and the previously aligned latent variables into GRU and update the network with the new alignment losses. This iterative process continues until all timestamps’ trajectory representations have been processed, enabling the network to grasp generalized trajectory features across different domains, thereby aligning the gap between domains. Furthermore, the pre-aligned representation module focuses on accounting for the interactions among trajectories at each timestamp for more comprehensive alignment, achieved through stepwise attention mechanisms. Social interactions along with the target trajectory features are regarded as trajectory representations at each timestamp, which are used for the recurrent alignment. Extensive experiments on three widely used benchmarks, including the ETH-UCY dataset [17, 27], SDD [29], and NBA dataset [43], show that our method outperforms all existing methods and demonstrate our superior generalization capabilities. The main contributions of our work are summarized as three-fold:
  • We study a brand new setting, named generalized pedestrian trajectory prediction. We perform a rigorous study of existing state-of-the-art methods and establish strong baselines in the setting of domain generalization.
@@
  • We present a recurrent alignment module to minimize the domain gap at both time-state and time-sequence levels. Furthermore, we introduce a pre-aligned representation module, which considers social interactions instead of solely target trajectories, into the domain alignment process.
2 Related Works
2.1 Pedestrian Trajectory Prediction
The objective of pedestrian trajectory prediction [49, 25, 5] is to anticipate future movements of pedestrians based on the observed trajectories and their surrounding context. This task faces several challenges, including human interactions [31, 20, 2], multimodality [30, 40, 3], long-tail distribution [38], and momentary prediction [34]. To tackle human interaction, SGCN [31] introduces a sparse graph convolution network designed to model sparse directed interactions and motion tendencies. GP-Graph [2] considers both individual-specific and group-based relationships, representing them as graph structures. In terms of multimodality, SIT [30] proposes a social interpretable tree to represent various possible future trajectories. SocialVAE [40] employs a conditional variational autoencoder (CVAE) to capture the distribution of multimodal future trajectories within a latent space. To handle the long-tail distribution problem, FEND [38] introduces an improved contrastive learning framework to effectively identify less frequent trajectory patterns. For momentary prediction, MOE [34] incorporates a momentary observation feature extractor, allowing efficient use of valuable momentary information in the prediction process. The recent work, T-GNN [41], has identified the domain shift problem and introduced a domain adaptation setting for pedestrian trajectory prediction. In this paper, we concentrate on a more practical setting, i.e., domain generalization, to tackle the domain shift problem for pedestrian trajectory prediction.
2.2 Domain Generalization
Domain generalization (DG) [13, 44, 36, 19] aims to acquire representations that remain invariant across various domains, enabling the model to effectively generalize to unfamiliar domains. Both domain generalization and domain adaptation (DA) [47, 37, 14, 24] tackle the challenges posed by domain shift. However, the most significant difference between DA and DG is that DG does not require access to the target domain during training. In real-world scenarios, DG is more practical and challenging than DA, mainly because collecting data from all target domains can be arduous in real-world situations. A previous work on pedestrian trajectory prediction [41] addresses the domain shift problem within the domain adaptation setting, which may be impractical in real-world scenarios. In a recent work, Frenet [42] solves the domain shift problem from the perspective of domain generalization. Nevertheless, it is important to note that this work focuses primarily on vehicle trajectory prediction, limiting its direct applicability to pedestrian trajectory prediction due to distinctions in scenarios, experimental settings, and technology frameworks. DG methods encompass various methods such as domain alignment [18, 4], meta-learning [36, 9], and data augmentation [35, 48]. Among these methods, domain alignment [18, 4] stands out as a widely used and effective strategy in DG, with the aim of minimizing the difference between the source domains to learn the domain-invariant representations. The rationale behind domain alignment is that features invariant to the source domain shift should also be robust to any unseen target domain shift. However, previous domain alignment methods are not tailored for sequence prediction tasks such as pedestrian trajectory prediction, as they do not consider the time sequence and human interaction properties. In this paper, we enhance the domain alignment method by incorporating a recurrent alignment strategy, tailoring it to be applicable for genralized pedestrian trajectory prediction.
@@ 2.2 Domain Generalization
3 Our Method
@@
3.1 Problem Formulation
The objective of pedestrian trajectory prediction is to predict the future trajectory coordinates of the agent based on its observed trajectory and neighboring trajectories. Mathematically, given a pedestrian in domain $i$, let $\mathbf{O}^{i}=[o_{i,1},o_{i,2},...,o_{i,T_{obs}}]$ be the observed trajectory over $T_{obs}$ timestamps, where $o_{i,t}\in\mathbb{R}^{2}$ denotes the 2D spatial coordinate at timestamp $t$ in domain $i$. Let $\mathcal{B}^{i}=[B_{i,1},B_{i,2},...,B_{i,T_{obs}}]$ be neighboring trajectories of $\mathbf{O}^{i}$, where $\mathbf{B}_{i,t}\in\mathbb{R}^{2\times a}$ is the 2D spatial coordinates of all neighboring trajectories at timestamp $t$. $a$ is the number of neighbors. The future trajectory is $\mathbf{Y}^{i}=[y_{i,T_{obs}+1},y_{i,T_{obs}+2},...,y_{i,T_{pred}}]$, where $y_{i,t}\in\mathbb{R}^{2}$ denotes the 2D spatial coordinate at timestamp $t$ in domain $i$. The goal is to train a prediction model $g(\cdot)$, so that the prediction $(\hat{\mathbf{Y}}^{i})^{min}$ with minimum error in multimodal predictions $(\hat{\mathbf{Y}}^{i})^{K}=g(\mathbf{O}^{i},\mathcal{B}^{i})$ is as close to the ground truth $\mathbf{Y}^{i}$ as possible.
@@ 3.1 Problem Formulation
3.2 Pre-Aligned Representation
Pedestrian trajectory prediction is more challenging than traditional time-sequence prediction tasks, because it necessitates accounting for both the time-sequence property and the human interaction property at the time-state level. To capture the trajectory interactions within a scene at each timestamp, we adopt a stepwise attention mechanism to model the influence of neighboring trajectories on the agent to be predicted. Given a trajectory $\mathbf{O}^{i}$ in domain $i$ and its neighboring trajectories $\mathcal{B}^{i}$, we first extract their high-dimensional feature embeddings at each timestamp as follows:
$\begin{split}\mathbf{F}_{i,t}^{o}=f_{o}(o_{i,t}),\quad\mathbf{F}_{i,t}^{b}=f_{b}(B_{i,t}),\\ \end{split}$ (1)
where $\mathbf{F}_{i,t}^{o}$ and $\mathbf{F}_{i,t}^{b}$ are the feature embeddings of $o_{i,t}$ and $B_{i,t}$, respectively. $f_{o}(\cdot)$ and $f_{b}(\cdot)$ are nonlinear multilayer perceptron (MLP) encoders. To consider the influence of neighboring agents, we adopt the attention mechanism to model the social interaction at each timestamp as follows:
@@ 3.2 Pre-Aligned Representation
where $\phi(\cdot,\cdot)$ denotes the linear transformation. $W_{i,t}^{q}$, $W_{i,t}^{k}$ and $W_{i,t}^{v}$ are learnable parameters. $\mathbf{Q}_{i,t}\in\mathbb{R}^{1\times d}$, $\mathbf{K}_{i,t}\in\mathbb{R}^{a\times d}$, and $\mathbf{V}_{i,t}\in\mathbb{R}^{a\times d}$ are the query, key and value of the attention at timestamp $t$ in domain $i$, respectively. $\mathbf{A}_{i,t}\in\mathbb{R}^{1\times d}$ is the output of the stepwise attention. $\sqrt{d_{k}}$ is a scaled factor to ensure numerical stability. Hence, we can formulate the trajectory representation at timestamp $t$ in domain $i$ by considering the observation $o_{i,t}$ and the stepwise attention $A_{i,t}$ as follows:
@@
where $X_{i,t}$ is the time-state representation of the trajectory to be predicted at timestamp $t$ in domain $i$. $\text{Concat}(\cdot)$ denotes the concatenation operation. Then, the representation of all trajectories in an arbitrary domain $i$ can be constructed as follows:
@@
where $F_{i}$ is the feature space of all trajectories in domain $i$. ${N_{i}}$ is the number of pedestrians in domain $i$. $\mathbf{f}_{i}^{(n)}$ is the trajectory’s pre-aligned representation of the $n^{th}$ pedestrian over $T_{obs}$ timestamps in domain $i$. $X_{i,t}^{n}$ is the time-state representation of the $n^{th}$ pedestrian’s trajectory at timestamp $t$ in domain $i$.
3.3 Recurrent Alignment
With the feature spaces $[F_{1},F_{2}]$ of two source domains not aligned properly, we aim to align them by minimizing the discrepancy of the two source domains. Then, we can assume that the resulting features will be domain-invariant and generalized well on the unseen target domains.

### v2

@@ Abstract
Pedestrian trajectory prediction is a crucial component in computer vision and robotics, but remains challenging due to the domain shift problem. Previous studies have tried to tackle this problem by leveraging a portion of trajectory data from the target domain to fine-tune the model. However, such domain adaptation methods are impractical in real-world scenarios, as it is infeasible to collect trajectory data from all potential target domains. In this paper, we study a new task named generalized pedestrian trajectory prediction, with the aim of generalizing the model to unseen domains without accessing their trajectories. To tackle this task, we further introduce a Recurrent Aligned Network (RAN) to minimize the domain gap through domain alignment. Specifically, we devise a recurrent alignment module to effectively align the trajectory feature spaces at both time-state and time-sequence levels by the recurrent alignment strategy. Furthermore, we introduce a pre-aligned representation module to combine social interactions with the recurrent alignment strategy, which aims to consider social interactions during the alignment process instead of just target trajectories. We extensively evaluate our method and compare it with state-of-the-art methods on three widely used benchmarks. The experimental results demonstrate the superior generalization capability of our method. Our work not only fills the gap in the generalization setting for practical pedestrian trajectory prediction, but also sets strong baselines in this field.
Index Terms:
Pedestrian Trajectory Prediction, Generalized Pedestrian Trajectory Prediction, Domain Generalization
I Introduction
@@
Pedestrian trajectory prediction aims to estimate the future movements of pedestrians given their observed trajectories. This task is crucial in various applications such as autonomous driving [2, 3, 4], visual recognition [5, 6, 7], navigation [8, 9, 10], and video surveillance [11, 12, 13]. Despite significant progress in this field, pedestrian trajectory prediction remains challenging due to the domain shift problem. This problem arises when a model encounters scenes significantly different from its training dataset, resulting in performance degradation. For example, the motion styles of students on a university campus differ significantly from those of players on a basketball court. Consequently, a predictor trained on campus datasets (e.g., SDD) may excel in that scene, but encounter difficulties with basketball court datasets (e.g., NBA), as depicted in Figure 1(a). Recent studies [14, 15] have identified the domain shift problem in pedestrian trajectory prediction and mitigated it through domain adaptation. However, in real-world scenarios like autonomous driving, domain changing is inevitable. Collecting trajectory data for all possible target domains is impractical, making the application of domain adaptation using data from target domains for model fine-tuning unfeasible in practice. Therefore, researching pedestrian trajectory prediction in more general scenarios, specifically under the domain generalization setting, is crucial for applications like autonomous driving.
In this paper, we tackle the challenge of predicting pedestrian trajectories in an unseen target domain dataset, without requiring data from the target domain during model training, as illustrated in Figure 1(b). We term this task as Generalized Pedestrian Trajectory Prediction. Generalized pedestrian trajectory prediction is more challenging and meaningful compared to previous settings. Specifically, unlike conventional settings that do not account for the domain shift problem, our approach is more suitable for the dynamically changing environments of autonomous driving. Compared to the previous domain adaptation setting, our method better addresses the difficulty of data collection in unknown scenarios.
One common and effective method to address this challenge is through domain alignment, which minimizes the data distribution gap between the source and target domains in high-dimensional feature spaces by using alignment losses such as Maximum Mean Discrepancy (MMD) [16], Correlation Alignment Distance (CORAL) [17], and adversarial loss [18]. Nevertheless, traditional domain alignment strategies are not well suited for the pedestrian trajectory prediction task because 1) they overlook the sequential property of the pedestrian trajectory prediction task and align different domains directly, without considering the alignment at both time-state and time-sequence levels within trajectories; 2) they overlook the surrounding alignment, ignoring that the predicted trajectory is not only associated with that specific trajectory but also related to other interconnected trajectories within the scene.
To address these issues, we further introduce a Recurrent Aligned Network (RAN), a new alignment framework tailored for generalized pedestrian trajectory prediction, as illustrated in Figure 2. RAN deals with domain alignment by considering both time-state and time-sequence alignments, along with the social interaction dynamics inherent in pedestrian trajectories. This advancement is facilitated by two key modules: a pre-aligned representation module and a recurrent alignment module. The recurrent alignment module aims to minimize the domain gap at both time-state and time-sequence levels by a novel recurrent alignment strategy. Specifically, we first input trajectory representations from different domains at the starting time step into the GRU units and update the network using alignment losses calculated by the output latent variables. Then, we input the trajectory representations of the next time step and the previously aligned latent variables into GRU and update the network with the new alignment losses. This iterative process continues until all time steps’ trajectory representations have been processed, enabling the network to grasp generalized trajectory features across different domains, thereby aligning the gap between domains. Furthermore, the pre-aligned representation module focuses on accounting for the human interactions at each time step for more comprehensive alignment, achieved through stepwise attention mechanisms. Specifically, social interactions along with the target trajectory features are regarded as trajectory representations at each time step, which are used for the recurrent alignment process. Finally, in addition to these two key modules, we adopt an RNN-based trajectory decoder to convert generalized trajectory features into multiple possible future trajectory predictions.
The main contributions of our work are summarized as three-fold:
  • We identify the domain shift problem and focus on a generalized approach for the pedestrian trajectory prediction task. Furthermore, we emphasize the differences in domain generalization between trajectory prediction and most other tasks, highlighting that domain generalization in trajectory prediction must consider its temporal and recurrent nature.
@@
  • We present a recurrent alignment module to minimize the domain gap at both time-state and time-sequence levels. Furthermore, we introduce a pre-aligned representation module, which considers human social interactions instead of solely target trajectories, into the domain alignment process.
  • Extensive experiments on three widely used benchmarks, including ETH-UCY, SDD, and NBA datasets, show that our method outperforms all existing methods and demonstrates our superior generalization capabilities.
II Related Work
TABLE I: Examples of domain gaps in different attributes between ETH-UCY, SDD and NBA datasets.
     Dataset      Main Scenario Type      Agent Type      # of Pedestrian      # of Trajectory
     ETH-UCY      Outdoor pedestrian paths      Pedestrian      1536      37270
     SDD      University campus            Bicyclist, Pedestrian,      Skateboarder, Cart, Car, Bus      5232      11323
     NBA      Basketball courts      Basketball player      3155      257230
II-A Pedestrian Trajectory Prediction
The objective of pedestrian trajectory prediction [19, 20, 21] is to anticipate future movements of pedestrians based on the observed trajectories and their surrounding context. This task faces several challenges, including human interactions [22, 23, 24], multimodality [25, 1, 26], long-tail distribution [27], and momentary prediction [28].
Taking advantage of advancements in deep learning, pedestrian trajectory prediction has made significant progress in addressing these challenges. For the challenge of human interaction, DifTraj [29] considers human intention features to enhance interaction modeling. MSRL [30] proposes to model complex human interactions using three separate branches: the temporal branch, the spatial branch, and the joint branch. Furthermore, due to the close relationship between tracking and trajectory prediction tasks, some interaction modeling methods [31, 32, 33, 34, 35] used in tracking tasks can also provide significant insights for trajectory prediction. In terms of multimodality challenge, SIT [25] proposes a social interpretable tree to represent various possible future trajectories and utilizes a teacher-student network to ensure that the learned future trajectory distributions align with the tree structure. SocialVAE [1] employs a conditional variational autoencoder (CVAE) to capture the distribution of multimodal future trajectories within a latent space. SICNet [36] proposes a method that first generates representative sparse instances for future modalities and then generates multimodal predictions conditioned on these sparse instances. To handle the long-tail distribution problem, FEND [27] introduces an improved contrastive learning framework to effectively identify less frequent trajectory patterns. For momentary prediction, MOE [28] incorporates a momentary observation feature extractor, allowing efficient use of valuable momentary information in the prediction process. Furthermore, the recent work, T-GNN [14], has identified the domain shift problem in pedestrian trajectory prediction, where a trained trajectory prediction model experiences performance degradation when transferred to unfamiliar environments. T-GNN addresses this issue using domain adaptation, where a portion of data from the target environment is used to fine-tune the model. However, this approach is not practical in trajectory prediction due to the difficulty in obtaining data from the target environment.
Thus, in this paper, we concentrate on a more practical setting, i.e., domain generalization, for pedestrian trajectory prediction. We also propose an efficient recurrent aligned neural network to tackle this domain shift problem under the domain generalization setting.
II-B RNNs for Sequence Prediction
Recurrent Neural Networks (RNNs) [37, 38, 39] constitute a versatile class of dynamic models that extend the capabilities of feedforward networks, enabling sequence generation in diverse tasks such as speech recognition [40], machine translation [41], and image captioning [42]. Pedestrian trajectory prediction is inherently a time-sequence prediction task, making it well-suited for feature extraction using RNN-based architectures like Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU). For example, PCCSNet [43] adopts the LSTM as encoders to extract observed and future trajectory features for deep clustering, classification, and further synthesis, and then generate the future prediction by the obtained features from the LSTM. Bitrap [44] extends the RNN into a GRU-based bi-directional decoder to improve the accuracy of long-term trajectory prediction and reduce collision rates. SocialVAE [1] introduces a temporal variational encoder architecture for pedestrian trajectory prediction, which combines RNNs with variational autoencoders (VAEs) at each time step. This integration incorporates multimodality into each time step, enhancing both the interpretability and accuracy of multimodal pedestrian trajectory prediction. However, pedestrian trajectory prediction is more complex than traditional time-sequence prediction, as the constantly changing environments make model generalization crucial. Traditional RNNs lack the ability to extract generalized features when handling temporal data across different scenes. As a result, these models may perform well in-domain but struggle with out-of-domain scenarios.
In this paper, we focus on addressing the domain generalization challenge in pedestrian trajectory prediction. To achieve this, we enhance the RNN with a time-sequential alignment strategy that employs recurrent alignment losses, effectively bridging the data gap across various domains.
@@ II-B RNNs for Sequence Prediction
II-C Domain Generalization
Domain generalization (DG) [45, 46, 47, 48] aims to acquire representations that remain invariant across various domains, enabling the model to effectively generalize to unfamiliar domains. Both domain generalization and domain adaptation (DA) [49, 50, 51, 52] tackle the challenges posed by domain shift. However, the most significant difference between DA and DG is that DG does not require access to the target domain during training. In real-world scenarios, DG is more practical and challenging than DA, mainly because collecting data from all target domains can be arduous in real-world situations. A recent work [14] demonstrates that different scenarios for pedestrian trajectory prediction have a domain shift problem, as illustrated in Table I, and addresses this domain shift problem within the domain adaptation setting, which may be impractical in real-world scenarios due to the difficulty of obtaining target domain data. Moreover, Frenet [53] solves the domain shift problem from the perspective of domain generalization. Nevertheless, it is important to note that this work focuses primarily on vehicle trajectory prediction, limiting its direct applicability to pedestrian trajectory prediction due to distinctions in scenarios, experimental settings, and technology frameworks.
Domain generalization methods encompass various methods such as domain alignment [54, 55], meta-learning [47, 56], and data augmentation [57, 58]. For example, AdapTraj [59] contributes to pedestrian trajectory prediction in generalization settings by specifically focusing on decoupling domain-invariant and domain-specific features of trajectories. MetaTra [60] solves the domain shift problem in pedestrian trajectory prediction through meta-learning. Unlike previous works, this paper places greater emphasis on the recurrent nature of trajectory data in domain generalization. Consequently, we have adopted a domain alignment strategy and enhanced it by incorporating a recurrent alignment strategy, tailoring it specifically for generalized pedestrian trajectory prediction.
III Proposed Method
@@
III-A Problem Formulation
The objective of pedestrian trajectory prediction is to predict the future trajectory coordinates of the agent based on its observed trajectory and neighboring trajectories. Mathematically, given a pedestrian in domain $i$, let $\mathbf{O}^{i}=[o_{i,1},o_{i,2},...,o_{i,T_{obs}}]$ be the observed trajectory over $T_{obs}$ time steps, where $o_{i,t}\in\mathbb{R}^{2}$ denotes the 2D spatial coordinate at time step $t$ in domain $i$. Let $\mathcal{B}^{i}=[B_{i,1},B_{i,2},...,B_{i,T_{obs}}]$ be the neighboring trajectories of $\mathbf{O}^{i}$, where $\mathbf{B}_{i,t}\in\mathbb{R}^{2\times a}$ is the 2D spatial coordinates of all neighboring trajectories at time step $t$. $a$ is the number of neighbors. The future trajectory is $\mathbf{Y}^{i}=[y_{i,T_{obs}+1},y_{i,T_{obs}+2},...,y_{i,T_{pred}}]$, where $y_{i,t}\in\mathbb{R}^{2}$ denotes the 2D spatial coordinate at time step $t$ in domain $i$. The goal is to train a prediction model $g(\cdot)$, so that the prediction $(\hat{\mathbf{Y}}^{i})^{min}$ with minimum error in multimodal predictions $(\hat{\mathbf{Y}}^{i})^{K}=g(\mathbf{O}^{i},\mathcal{B}^{i})$ is as close to the ground truth $\mathbf{Y}^{i}$ as possible.
@@ III-A Problem Formulation
III-B Pre-Aligned Representation
Pedestrian trajectory prediction is more challenging than traditional time-sequence prediction tasks, because it necessitates accounting for both the time-sequence property and the human interaction property at the time-state level. Before modelling the time-sequence information of pedestrian trajectories, we first model the interactions between pedestrians at each time step. To capture the trajectory interactions within a scene at each time step, we adopt a stepwise attention mechanism to model the influence of neighboring trajectories on the agent to be predicted.
Given a trajectory $\mathbf{O}^{i}$ in domain $i$ and its neighboring trajectories $\mathcal{B}^{i}$, we first extract their high-dimensional feature embeddings at each time step as follows:
$\begin{split}\mathbf{F}_{i,t}^{o}&=f_{o}(o_{i,t}),\\ \quad\mathbf{F}_{i,t}^{b}&=f_{b}(B_{i,t}),\\ \end{split}$ (1)
where $\mathbf{F}_{i,t}^{o}$ and $\mathbf{F}_{i,t}^{b}$ are the feature embeddings of $o_{i,t}$ and $B_{i,t}$, respectively. $f_{o}(\cdot)$ and $f_{b}(\cdot)$ are nonlinear multilayer perceptron (MLP) encoders. To consider the influence of neighboring agents, we adopt the attention mechanism to model the social interaction at each time step as follows:
@@ III-B Pre-Aligned Representation
where $\phi(\cdot,\cdot)$ denotes the linear transformation. $W_{i,t}^{q}$, $W_{i,t}^{k}$ and $W_{i,t}^{v}$ are learnable parameters. $\mathbf{Q}_{i,t}\in\mathbb{R}^{1\times d}$, $\mathbf{K}_{i,t}\in\mathbb{R}^{a\times d}$, and $\mathbf{V}_{i,t}\in\mathbb{R}^{a\times d}$ are the query, key and value of the attention at time step $t$ in domain $i$, respectively. $\mathbf{A}_{i,t}\in\mathbb{R}^{1\times d}$ is the output of the stepwise attention. $\sqrt{d_{k}}$ is a scaled factor to ensure numerical stability. The feature of the target pedestrian at the time step $t$ should consist of its own coordinate information and its interaction with surrounding pedestrians. Hence, we can formulate the trajectory representation at time step $t$ in domain $i$ by considering the observation $o_{i,t}$ and the stepwise attention $A_{i,t}$ as follows:
@@
where $X_{i,t}$ is the time-state representation of the trajectory to be predicted at time step $t$ in domain $i$. $\text{Concat}(\cdot)$ denotes the concatenation operation.
Pedestrian trajectories consist of time-sequence information, which is constructed from time-state information. Hence, the representation of all trajectories in an arbitrary domain $i$ can be constructed by time-state representation $X_{i,t}$ as follows:
@@
where $F_{i}$ is the feature space of all trajectories in domain $i$. ${N_{i}}$ is the number of pedestrians in domain $i$. $\mathbf{f}_{i}^{(n)}$ is the trajectory’s pre-aligned representation of the $n^{th}$ pedestrian over $T_{obs}$ time steps in domain $i$. $X_{i,t}^{n}$ is the time-state representation of the $n^{th}$ pedestrian’s trajectory at time step $t$ in domain $i$.
III-C Recurrent Alignment
As previously discussed, the goal of the recurrent alignment module is to eliminate the domain gap by extracting generalized trajectory features across different scenarios. Hence, with the feature spaces $[F_{1},F_{2}]$ of two source domains not aligned properly, we can align them by minimizing the discrepancy of the two source domains. Then, we can assume that the resulting features will be domain-invariant and generalized well on the unseen target domains.
where $h_{i,t}^{n}$ is the hidden state of the trajectory’s time-state representation $X_{i,t}^{n}$. $r_{i,t}^{n}$, $z_{i,t}^{n}$, and $g_{i,t}^{n}$ are the reset, update, and new gates, respectively. $W$ and $b$ denote learning parameters. $\sigma$ is the activation function. $\odot$ is the element-wise multiplication. The initial hidden state $h_{i,0}$ is produced by a linear transformation of the feature embedding $\mathbf{F}_{i,t}^{b}$. Hence, the feature space of each domain at different time steps can be represented as follows:
$\begin{split}S_{i,t}=\sum_{n=1}^{N_{i}}h_{i,t}^{n},\end{split}$ (6)
where $S_{i,t}$ is the feature space of all trajectories in domain $i$ at time step $t$. Given $S_{i,t}$, we can minimize the discrepancy between the two source domains using the proposed recurrent alignment strategy and the recurrent alignment losses outlined in Algorithm 1. $\mathcal{L}_{rec}(t)$ is the recurrent alignment loss $\mathcal{L}_{rec}$ at time step $t$. Then, our network is capable of modeling generalized trajectory features, which can be regarded as domain-invariant.